# ppa_t_p.bib

### v1

@@ 5 Anonymous Credentials: The Practice
It seems fair to say that when considering industry adoption, anonymous credentials have not been a success story so far. Fortunately, the picture is changing and we see increasing interest in deploying such technology. In this section we want to look at the anonymous credential technology from an implementation, standardization and deployment perspective.
@@ 5.1 Implementations
To the best of our knowledge, the first implementation of a multi-use anonymous credential system was the idemix system developed within IBM [DBLP:conf/ccs/CamenischH02] (using strong RSA-based CL signatures [DBLP:conf/scn/CamenischL02]) and further developed and improved among others over a series of EU projects such as the EU FP7 projects PrimeLife and ABC4Trust. Later, the idemix project has been integrated in the Hyperledger project and is the basis for the Hyperledger AnonCreds [^6] https://lf-hyperledger.atlassian.net/wiki/spaces/ANONCREDS/overview and related projects (e.g., Aries Bifold or Hyperledger Aries), which plan to move to BBS signatures in its version 2.0 [BeKaBBS23]. Moreover, the first implementation of a single-use anonymous credential system relying on the techniques in [BrandsBook] was due to Brands at Credentica and was then acquired by and continued as the U-Prove project at Microsoft [^7] https://www.microsoft.com/en-us/research/project/u-prove/.
We note that in the field of distributed ledger technology and verifiable credentials, there are meanwhile numerous open-source implementations of anonymous credentials based on BBS(+) signatures. We just want to mention the Rust bbs crate [^8] https://crates.io/crates/bbs, the docknetwork implementation [^9] https://github.com/docknetwork/crypto and the BBS implementation of Digital Bazaar [^10] https://github.com/digitalbazaar/bbs-signatures here for some (random) examples.
The anonymous credentials zoo [^11] https://tokenzoo.github.io provides an (somewhat outdated) overview of some multi- as well as single-show anonymous credentials with some pointers to implementations. Unfortunately, we are not aware of a place that compiles a list of current open-source implementations, something that would be very helpful and highly desirable.
@@ 5.2 Standardization Efforts
First we want to mention the Crypto Forum Research Group (CFRG) within the Internet Engineering Task Force (IETF), which publishes Requests For Comments (RFCs). There are ongoing standardization processes for BBS signatures [I-D.irtf-cfrg-bbs-signatures], blind BBS signatures [I-D.kalos-bbs-blind-signatures] as well as pseudonyms for BBS signatures [I-D.kalos-bbs-per-verifier-linkability]. Moreover, there are many (ongoing) standardizations for related primitives such as oblivious pseudo-random functions (OPRFs) or blind signatures.
In addition, there are non-profit organizations such as the Linux Foundation [^12] https://www.linuxfoundation.org or industry consortia such as the Decentralized Identity Foundation (DIF) [^13] https://identity.foundation, the World-Wide Web Consortium (W3C) [^14] https://www.w3.org or ZKProof [^15] https://zkproof.org who are running working groups on anonymous credential technologies and the underlying technologies respectively.
Finally, there is the International Standardization Organization (ISO) currently running a preliminary work item (ISO/IEC PWI 24843) on Privacy-Preserving Attribute-Based Credentials within the ISO/IEC JTC1/SC 27 WG 2 and on guidelines on privacy preservation based on zero-knowledge proofs (ISO/IEC CD 27565). Moreover, the European Telecommunications Standards Institute (ETSI) has several technical reports on related technology, such as ETSI TR 119 476161616https://www.etsi.org/deliver/etsi_tr/119476 on selective disclosure and zero-knowledge proofs.
@@ 5.3 Real-World Deployments
The biggest real-world use-cases related to anonymous credential technology and rolled out on billions of devices are the Direct Anonymous Attestation (DAA) scheme implemented in any Trusted Platform Module (TMP) or the Enhanced Privacy ID (EPID) [^17] https://www.intel.com/intel-enhanced-privacy-id-epid-security-technology available in Intel’s SGX. However, while somewhat close to multi-use anonymous credentials, those schemes were designed for specific use-cases, do not consider attributes, and are thus not generally applicable.
@@
There are also various governmental use-cases that rely on multi-use anonymous credentials, e.g., supporting sustainable economic development by the government of British Columbia (Canada) or digital travel credentials by the government of Aruba. We refer to [cryptoeudi] for more discussion on such use-cases.
Apart from this we want to mention the use of so called anonymous tokens in various forms and use-cases. Those represent single-use credentials with or without attributes that are either privately or publicly available. This includes Privacy Pass [DBLP:journals/popets/DavidsonGSTV18] as used by Cloudflare [^20] https://blog.cloudflare.com/privacy-pass-standard/ or the Anonymous Credential Service by Meta [^21] https://engineering.fb.com/2022/12/12/security/anonymous-credential-service-acs-open-source/.
@@ 6 European Union Digital Identity Wallet (EUDIW)
With the EU Digital Identity Regulation entering into force in May 2024, the EU proposes the EU Digital Identity Wallet (EUDIW), which shall be a “fully mobile, secure and user-friendly” service, enabling users to identify themselves to public and private online services. Consequently, all member states are required to offer such an EUDI wallet to their citizens and residents by 2026. We refer to [cryptoeudi] for a more thorough discussion and will only briefly discuss the initial design presented in the architectural reference framework (ARF) version 1.4.0 [arf] and its shortcomings from a privacy perspective.
@@ 6.1 Signatures on Salted Hashes and their Privacy Issues
The approach pursued in the ARF version 1.4.0 [arf] is to follow the realization of selective attribute disclosure credentials from conventional digital signatures (e.g., ECDSA) as specified in the ISO/IEC 18013-5 (mobile driving license) standard. On a high-level the idea is that an issuer signs a list of “salted-hashes” $(h_{1},\ldots,h_{n})$ instead of directly signing the attributes $(a_{1},\ldots,a_{n})$. Each salted hash $h_{i}=H(a_{i}||r_{i})$ is computed using a cryptographic hash function $H$ such as SHA-3 on input the attribute $a_{i}$ concatenated with a sufficiently long uniformly random string $r_{i}$. This can be viewed as a hash commitment, i.e., $h_{i}$ hides $a_{i}$ as long as $r_{i}$ is not known and hash value $h_{i}$ cannot be opened to a different $a_{i}^{\prime}\neq a_{i}$ and thus is binding.
The idea is now that when a user presents a signature for $(h_{1},\ldots,h_{n})$ to some RP, for attributes it wants to show, it reveals $a_{i}$ along with $r_{i}$. For all attributes that are not disclosed, the user simply reveals $h_{j}$, which as discussed does not reveal anything about $a_{j}$ (as $r_{j}$ is not revealed).
Privacy discussion. This represents the functionality of selective disclosure, i.e., the user can decide which attributes $a_{i}$ to reveal. However, when showing the same signature multiple times, the signature as well as $(h_{1},\ldots,h_{n})$ is always revealed and thus can be trivially linked by any RP. This can be avoided by showing every signature only once, i.e., viewing them as single-use credentials, which comes at the cost of requiring frequent re-issuance of new credentials and keeping track of already used signatures on the user side. Moreover, the collusion of the IdP and any RP always completely breaks unlinkability as the IdP issues all signatures and thereby sees all attributes. We also refer the interested reader to [cryptoeudi] for a more thorough discussion of these aspects.
@@ 7 Barriers to Widespread Adoption
While there might be numerous reasons that slow down or prevent a widespread adoption of “new” technologies such as anonymous credentials, we will solely focus on technical aspects. And even here, we do not claim to be exhaustive. We will discuss two aspects which we consider really important (and non-trivial to solve) issues that complicate practical adoption. They have also shown up in a recent discussion about the EUDIW after providing the feedback in [cryptoeudi]. Moreover, at the end of this section, we will also briefly discuss challenges that are arising from the recently started transitions to post-quantum cryptography.
@@ 7.1 Dependency on Established Standards
Moreover, there might be standards that do not directly relate to the AC schemes and their technical details, that still can influence potential deployments of such schemes. One example is how credentials in such schemes are specified when it comes to the document structure. Here the two currently dominating standards are verifiable credentials by the W3C [^22] https://www.w3.org/TR/vc-overview/ as well as the mdoc format specified in the ISO/IEC 18013 family of documents that specify the technical and operational requirements for physical and mobile driver’s licenses (mDL). While in the design of the former the use of anonymous credentials has been considered, the latter does not do so. Actually, it is considered along with a dedicated solution to selective disclosure, which as we have discussed in Section 6.1 does not provide a sufficient degree of unlinkability. This might introduce overhead and the requirement for ad-hoc intermediate formats when using them along with anonymous credentials.
@@ 7.2 Hardware Support and Compatibility
While only requiring pure software implementations of new technologies such as anonymous credentials is a manageable task, the integration of new cryptographic approaches into existing ecosystems and infrastructure is a much more delicate task. In the former case it boils down to secure implementations of cryptography, which by itself is a non-trivial task and requires lots of expert knowledge. But it does not suffer from critical external dependencies. However, in the latter case it requires (backwards) compatibility at many places. For instance, cryptographic keys and operations involving secret keys (such as signing) are typically handled by dedicated hardware such as hardware security modules (HSM), Trusted Platform Modules (TPMs) or other secure elements such as smart cards. Alternatively, they can be handled by Trusted Execution Environments (TEEs) such as ARM’s TrustZone or Intel’s SGX. Irrespective of the concrete realization, what is common to all those technologies is that the provided cryptographic functionality is limited to very few cryptographic primitives. When it comes to signature functionality, this typically either means RSA or ECDSA and in the latter case specific standardized elliptic curves such as curve P-256 (aka secp256r1 or prime256v1).
However, anonymous credentials such as those based on BBS [DBLP:conf/crypto/BonehBS04], as suggested in [cryptoeudi], require specific pairing-friendly elliptic curves such as BN-462 or BLS12-381232323https://members.loria.fr/AGuillevic/pairing-friendly-curves/ that are not readily available in the aforementioned technologies. Moreover, this requires other interfaces as those available for conventional signatures. Consequently, such a functionality is simply not available in currently deployed secure elements and TEEs. The only functionality that comes close is that of Direct Anonymous Attestation (DAA) in TMPs or EPID available in Intel’s SGX. However, as already mentioned they were designed for the specific use-cases and do not consider attributes in any way. While making anonymous credentials available in TEEs seems easier than when they are rooted in hardware, such changes might also require significant changes to software stacks, e.g., client platforms such as the Android system. Especially as such systems also require backwards compatibility. A very informative recent talk on changing legacy cryptography and the associated problems [^24] https://youtu.be/-YXCojP8IjE by abhi shelat is highly recommended at this place.
@@ 7.3 Transitioning to Post-Quantum Cryptography
It is important to mention though, that while we now see the first NIST standards for basic public-key cryptographic primitives (i.e., public-key encryption and signatures), we are still not close to having a clear picture of candidates for post-quantum anonymous credentials. There are recent constructions of lattice-based anonymous credentials [DBLP:conf/crypto/BootleLNS23, DBLP:journals/iacr/ArgoGJLRS24] and first proof-of-concept implementations are available [cryptoeprint:2024/1846]. While these constructions do not yet seem well-studied enough in terms of security and efficiency, it can however be expected that significant research will go into this topic in the upcoming years. Moreover, as discussed in more detail in [cryptoeudi], post-quantum security is less critical in authentication primitives than it is for encryption. Also it should be noted that if the privacy property of an anonymous credential system is unconditional (as it is the case for the BBS family mentioned in Section 4.3), then even when instantiated from building blocks that could be broken by a hypothetical quantum computer, the privacy is not endangered and will hold forever (even when given unlimited computing power).
@@ 8 Conclusions and Outlook
It needs to be concluded, that currently enrolled (and partly also planned) identity solutions typically provide rather weak privacy protection and are quite far from what could be considered the ideal case. Anonymous credentials are the right tool to achieve strong privacy protection and are well studied and well understood by the research community. Unfortunately, although they have already been proposed back in the early 1980s by Chaum, and a large body of literature as well as practical use-cases and open-source implementations exist, for a long time they did not see adoption in industry. Fortunately, in recent years we see a growing interest from the industry and governments to adopt such privacy-preserving technologies. However, deploying new technology comes with huge (financial) efforts, particularly when use-cases require backwards compatibility with existing hardware implementing only “legacy cryptography”, which typically cannot be readily used with approaches and in particular modern anonymous credentials. Recent progress in cryptographic research and in parricular in the field of zk-SNARKs, however, does now even enable solutions based on existing “legacy cryptography” and so even immediate large-scale deployments do not seem out of reach.
Nevertheless, when deploying solutions now, they are likely here to stay for quite some time. Thus, solid and forward-looking planning is essential. This particularly holds true for the governmental domain and large scale solutions like the EUDIW. For instance, when considering to rolling out a new solution today, then it is definitely advisable to consider quantum resistance (i.e., post-quantum security).
At the time of writing this paper, the technical implementation of the EUDIW has been postponed [^26] https://epicenter.works/en/content/finally-a-no-to-overreaching-id-systems. At this point it is still unclear how the final version of the ARF will look like and which technology will ultimately be adopted. We strongly hope that it will be a solution that considers strong privacy protection.
Acknowledgments. The author is very grateful to René Mayrhofer, Omid Mir, Octavio Perez-Kempner and Mahdi Sedaghat for their helpful feedback on a draft of this paper.

### v2

@@ 5 Anonymous Credentials: The Practice
It is fair to say that when considering industry adoption, anonymous credentials have not been a success story so far. Fortunately, the picture is changing and we see increasing interest in deploying such technology. In this section we want to look at the anonymous credential technology from an implementation, standardization and deployment perspective.
@@ 5.1 Implementations
To the best of our knowledge, the first implementation of a multi-use anonymous credential system was the idemix system developed within IBM in the early 2000’s [DBLP:conf/ccs/CamenischH02] (using strong RSA-based CL signatures [DBLP:conf/scn/CamenischL02]) and further developed and improved among others over a series of EU projects such as the EU FP7 projects PrimeLife and ABC4Trust. Later, the idemix project has been integrated into the Hyperledger project, now sponsored and managed by the Linux Foundation [^8] https://www.linuxfoundation.org, and is the basis for the Hyperledger AnonCreds [^9] https://lf-hyperledger.atlassian.net/wiki/spaces/ANONCREDS/overview and related projects (e.g., Aries Bifold or Hyperledger Aries), which plan to move to BBS signatures in its version 2.0 [BeKaBBS23]. Moreover, the first implementation of a single-use anonymous credential system relying on the techniques in [BrandsBook] was due to Brands at Credentica in the early 2000’s and was then acquired by and continued as the U-Prove project at Microsoft [^10] https://www.microsoft.com/en-us/research/project/u-prove/.
We note that in the field of distributed ledger technology and verifiable credentials, there are meanwhile numerous open-source implementations of anonymous credentials based on BBS(+) signatures. We may mention here the Rust bbs [^11] https://crates.io/crates/bbs or ursa crate [^12] https://docs.rs/ursa/, the docknetwork implementation [^13] https://github.com/docknetwork/crypto and the BBS implementation of Digital Bazaar [^14] https://github.com/digitalbazaar/bbs-signatures, provided solely as some illustrative instances. We also refer the interested reader to a recent work by Flamini et al. [DBLP:journals/istr/FlaminiSSSTR24] for a comparison and experimental evaluation.
The anonymous credentials zoo [^15] https://tokenzoo.github.io provides an (somewhat outdated) overview of some multi- as well as single-use anonymous credentials with some pointers to implementations. Unfortunately, we are not aware of a place that compiles a list of current open-source implementations, something that would be very helpful and highly desirable.
@@ 5.2 Standardization Efforts
First, we want to mention the Crypto Forum Research Group (CFRG) within the Internet Engineering Task Force (IETF), an open, global engineering community that publishes de facto standards called Requests For Comments (RFCs). There are ongoing standardization processes for BBS signatures [I-D.irtf-cfrg-bbs-signatures], blind BBS signatures [I-D.kalos-bbs-blind-signatures] as well as pseudonyms for BBS signatures [I-D.kalos-bbs-per-verifier-linkability]. Moreover, there are many (ongoing) standardization efforts for related primitives such as oblivious pseudo-random functions (OPRFs) or blind signatures.
In addition, there are non-profit organizations such as the aforementioned Linux Foundation or industry consortia such as the Decentralized Identity Foundation (DIF) [^16] https://identity.foundation, the World Wide Web Consortium (W3C) [^17] https://www.w3.org or ZKProof [^18] https://zkproof.org which are running working groups on anonymous credential technologies and the underlying technologies respectively.
Finally, there is the International Organization for Standardization (ISO) currently running a preliminary work item (ISO/IEC PWI 24843) on Privacy-Preserving Attribute-Based Credentials within the ISO/IEC JTC1/SC 27 WG 2 and on guidelines on privacy preservation based on zero-knowledge proofs (ISO/IEC CD 27565). Moreover, the European Telecommunications Standards Institute (ETSI) has several technical reports on related technology, such as ETSI TR 119 476191919https://www.etsi.org/deliver/etsi_tr/119476 on selective disclosure and zero-knowledge proofs.
@@ 5.3 Real-World Deployments
The biggest real-world use cases related to anonymous credential technology and rolled out on billions of devices are the Direct Anonymous Attestation (DAA) scheme implemented in any Trusted Platform Module (TPM) or the Enhanced Privacy ID (EPID) [^20] https://www.intel.com/intel-enhanced-privacy-id-epid-security-technology available in Intel’s SGX. However, while somewhat close to multi-use anonymous credentials, those schemes were designed for specific use cases, do not consider attributes, and are thus not generally applicable.
@@
There are also various governmental use cases that rely on multi-use anonymous credentials, e.g., supporting sustainable economic development by the government of British Columbia (Canada) or digital travel credentials by the government of Aruba. We refer to [cryptoeudi] for more discussion on such use cases.
Apart from this, we want to mention the use of so-called anonymous tokens in various forms and use cases. Those represent single-use credentials, with or without attributes that are either privately or publicly available. This includes Privacy Pass [DBLP:journals/popets/DavidsonGSTV18] as used by Cloudflare [^23] https://blog.cloudflare.com/privacy-pass-standard/, or the Anonymous Credential Service by Meta [^24] https://engineering.fb.com/2022/12/12/security/anonymous-credential-service-acs-open-source/.
@@ 6 European Union Digital Identity Wallet (EUDIW)
With the EU Digital Identity Framework Regulation [^25] https://oeil.secure.europarl.europa.eu/oeil/en/procedure-file?reference=2021/0136(COD) (commonly known as eIDAS 2.0) entering into force in May 2024, the EU proposes the EU Digital Identity Wallet (EUDIW), which shall be a “fully mobile, secure and user-friendly” service, enabling users to identify themselves to public and private online services. Consequently, all member states are required to offer such an EUDIW to their citizens and residents by 2026. We refer to [cryptoeudi] for a more thorough discussion, and will only briefly discuss the initial design presented in the architectural reference framework (ARF) version 1.4.0 [arf] and its shortcomings from a privacy perspective.
@@ 6.1 Signatures on Salted Hashes and their Privacy Issues
The approach pursued in the ARF version 1.4.0 [arf] is to follow the realization of selective attribute disclosure credentials from conventional digital signatures (e.g., ECDSA) as specified in the ISO/IEC 18013-5 (mobile driving license) standard. At a high level, the idea is that an issuer signs a list of “salted hashes” $(h_{1},\ldots,h_{n})$ instead of directly signing the attributes $(a_{1},\ldots,a_{n})$. Each salted hash $h_{i}=H(a_{i}||r_{i})$ is computed using a cryptographic hash function $H$ such as SHA-3 on input the attribute $a_{i}$ concatenated with a sufficiently long uniformly random string $r_{i}$. This can be viewed as a hash commitment, i.e., $h_{i}$ hides $a_{i}$ as long as $r_{i}$ is not known and hash value $h_{i}$ cannot be opened to a different $a_{i}^{\prime}\neq a_{i}$, and thus is binding. Consequently, the commitment hides the attribute until the salt is revealed; even if the same attribute appears in different credentials, the added salt ensures that each hash is unique, thereby preventing straightforward linkage.
The idea is now that when a user presents a signature for $(h_{1},\ldots,h_{n})$ to some RP for attributes it wants to present, it reveals $a_{i}$ along with $r_{i}$. For all attributes that are not disclosed, the user simply reveals $h_{j}$, which as discussed does not reveal anything about $a_{j}$ (as $r_{j}$ is not revealed).
Privacy discussion. This represents the functionality of selective disclosure, i.e., the user can decide which attributes $a_{i}$ to reveal. However, when presenting the same signature multiple times, the signature as well as $(h_{1},\ldots,h_{n})$ are always revealed and thus can be trivially linked by any RP. This can be avoided by presenting every signature only once, i.e., viewing them as single-use credentials, which comes at the cost of requiring frequent re-issuance of new credentials and keeping track of already used signatures on the user side, and may reduce convenience. Moreover, the collusion of the IdP and any RP always completely breaks unlinkability as the IdP issues all signatures and thereby sees all attributes. We also refer the interested reader to [cryptoeudi] for a more thorough discussion of these aspects.
@@ 7 Barriers to Widespread Adoption
While there might be numerous reasons that slow down or prevent a widespread adoption of “new” technologies such as anonymous credentials, we will solely focus on technical aspects. And even here, we do not claim to be exhaustive and do not cover important aspects such as ensuring end-user usability or system interoperability. We will discuss two aspects which we consider really important (and non-trivial to solve) issues that complicate practical adoption. They have also shown up in a recent discussion about the EUDIW after providing the feedback in [cryptoeudi]. Moreover, at the end of this section, we will also briefly discuss challenges that are arising from the recently started transitions to post-quantum cryptography.
@@ 7.1 Dependency on Established Standards
Moreover, there might be standards that do not directly relate to the AC schemes and their technical details, that still can influence potential deployments of such schemes. One example is how credentials in such schemes are specified when it comes to the document structure. Here, the two currently dominating standards are verifiable credentials by the W3C [^26] https://www.w3.org/TR/vc-overview/ and the mdoc format specified in the ISO/IEC 18013 family of documents that specify the technical and operational requirements for physical and mobile driver’s licenses (mDL). While in the design of the former the use of anonymous credentials has been considered, the latter does not do so. Actually, it is considered along with a dedicated solution to selective disclosure, which as we have discussed in Section 6.1 does not provide a sufficient degree of unlinkability. This might introduce overhead and the requirement for ad-hoc intermediate formats when using them along with anonymous credentials.
@@ 7.2 Hardware Support and Compatibility
While only requiring pure software implementations of new technologies such as anonymous credentials is a manageable task, the integration of new cryptographic approaches into existing ecosystems and infrastructure is a much more delicate task. In the former case it boils down to secure implementations of cryptography, which by itself is a non-trivial task and requires lots of expert knowledge. But it does not suffer from critical external dependencies. However, in the latter case it requires backward compatibility at many places. For instance, cryptographic keys and operations involving secret keys (such as signing) are typically handled by dedicated hardware such as hardware security modules (HSM), Trusted Platform Modules (TPMs), or other secure elements such as smart cards. Alternatively, they can be handled by Trusted Execution Environments (TEEs) such as ARM’s TrustZone or Intel’s SGX. Irrespective of the concrete realization, what is common to all those technologies is that the provided cryptographic functionality is limited to very few cryptographic primitives. When it comes to signature functionality, this typically either means RSA or ECDSA, and, in the latter case, specific standardized elliptic curves such as curve P-256 (also known as secp256r1 or prime256v1).
However, anonymous credentials such as those based on BBS [DBLP:conf/crypto/BonehBS04], as suggested in [cryptoeudi], require specific pairing-friendly elliptic curves such as BN-462 or BLS12-381272727https://members.loria.fr/AGuillevic/pairing-friendly-curves/ that are not readily available in the aforementioned technologies. Moreover, this requires other interfaces as those available for conventional signatures. Consequently, such a functionality is simply not available in currently deployed secure elements and TEEs. The only functionality that comes close is that of Direct Anonymous Attestation (DAA) in TPMs or EPID available in Intel’s SGX. However, as already mentioned they were designed for the specific use cases and do not consider attributes in any way. While making anonymous credentials available in TEEs seems easier than when they are rooted in hardware, such changes might also require significant changes to software stacks, e.g., client platforms such as the Android system. Especially as such systems also require backward compatibility. A very informative recent talk on changing legacy cryptography and the associated problems [^28] https://youtu.be/-YXCojP8IjE by abhi shelat is highly recommended at this place.
@@ 7.3 Transitioning to Post-Quantum Cryptography
It is important to mention though, that while we now see the first NIST standards for basic public-key cryptographic primitives (i.e., public-key encryption and signatures), we are still not close to having a clear picture of candidates for post-quantum anonymous credentials. There are recent constructions of lattice-based anonymous credentials [DBLP:conf/crypto/BootleLNS23, DBLP:journals/iacr/ArgoGJLRS24, cryptoeprint:2025/356] and first proof-of-concept implementations are available from IBM [cryptoeprint:2024/1846] or the EU QUBIP project [^30] https://github.com/Cybersecurity-LINKS/pqzk-blns. While these constructions do not yet seem well-studied enough in terms of security and efficiency, it can, however, be expected that significant research will go into this topic in the upcoming years. Moreover, as discussed in more detail in [cryptoeudi], post-quantum security is less critical in authentication primitives than it is for encryption. Also it should be noted that if the privacy property of an anonymous credential system is unconditional (as it is the case for the BBS family mentioned in Section 4.4), then even when instantiated from building blocks that could be broken by a hypothetical quantum computer, the privacy is not endangered and will hold forever (even when given unlimited computing power).
@@ 8 Conclusions and Outlook
It needs to be concluded that currently enrolled (and partly also planned) identity solutions typically provide rather weak privacy protection and are quite far from what could be considered the ideal case. Anonymous credentials are the right tool to achieve strong privacy protection and are well studied and well understood by the research community. However, although they have already been proposed back in the early 1980s by Chaum, and a large body of literature as well as practical use cases and open-source implementations exist, for a long time they did not see adoption in industry. Fortunately, in recent years we see a growing interest from the industry and governments to adopt such privacy-preserving technologies. However, deploying new technology comes with huge (financial) efforts, particularly when use cases require backward compatibility with existing hardware that only implements “legacy cryptography”, which typically cannot be readily used with approaches and in particular modern anonymous credentials. Recent progress in cryptographic research and in particular in the field of zk-SNARKs, however, does now even enable solutions based on existing “legacy cryptography” and so even immediate large-scale deployments do not seem out of reach.
Nevertheless, when deploying solutions now, they are likely here to stay for quite some time. Thus, solid and forward-looking planning is essential. This particularly holds true for the governmental domain and large scale solutions like the EUDIW. For instance, when considering the rollout a new solution today, it is definitely advisable to consider quantum resistance (i.e., post-quantum security).
At the time of writing this paper, the technical implementation of the EUDIW is still under discussion. [^31] https://github.com/eu-digital-identity-wallet/eudi-doc-architecture-and-reference-framework/discussions At this point it cannot be determined with certainty how the final version of the ARF will look like and which technology will ultimately be adopted. We strongly hope that it will be a solution that considers strong privacy protection.
Acknowledgments. The author is very grateful to Yod Samuel Martín, René Mayrhofer, Omid Mir, Octavio Perez-Kempner and Mahdi Sedaghat for their helpful feedback on a draft of this paper.